OpenStudy (anonymous):

Find a model for simple harmonic motion if the position at t = 0 is 2, the amplitude is 2 centimeters, and the period is 6 seconds.

OpenStudy (anonymous):

x=Acosωt where A=2 cm ω=2\(\pi\)/6 rad/s You can see that at t=0 ,x=2 cm

OpenStudy (souvik):

x=Asin\(\omega\)(t+\(\alpha\)) t=0 x=2 cm A=2cm \(\omega=2\pi/6\) rad/s so...\(\alpha\)=90 degree so the equation is x=Asin\((2\pi/6*t+\pi/2)\) or x=Acos\(2\pi/6*t\)